Circular No. 6193

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

(2060) CHIRON = COMET 95P/CHIRON M. Womack, Pennsylvania State University, Erie; and S. A. Stern,

Southwest Research Institute, report: “We obtained a first detection of CO in Chiron via the molecule’s J=1–0 rotational transition at 115 GHz (2.6 mm), during June 10–12 UT with the National Radio Astronomy Observatory 12-m telescope on Kitt Peak. The line was observed in two polarizations, with each of two independent filterbanks of 100- and 250-kHz spectral resolution. The line was blueshifted by ∼ 0.5 km/s from the ephemeris velocity. The CO line area was measured to be 0.013 ± 0.005 K km/s in both sets of data (at 5σ in the 100-kHz spectrum and 3σ in the 250-kHz data). Assuming that the CO fills the 50′′ NRAO beam and that the excitation, rotational, and kinetic temperatures of the gas are all between 10 and 50 K, we derive a CO column density of (2 ± 1) × 10^{13} cm^{−2}. A Haser model calculation, assuming an isotropic coma expansion at velocities of 0.5 km/s, implies a production rate of Q(CO) = (4 ± 2) × 10^{28} mol s^{−1} if CO is a parent molecule. Additional observations of CO in Chiron’s coma during the upcoming perihelion passage are encouraged.”

QZ VULPECULAE P. A. Charles and J. Casares, Oxford University, write: “Optical spec-

troscopy of QZ Vul obtained with the 4.2-m William Herschel Telescope on La Palma shows broad (FWHM 2000 km/s), double-peaked Hα emission superposed on the late-type spectrum of the secondary star, detected here for the first time and constrained to be in the range K3–K7; QZ Vul is the quiescent (mag 21) counterpart of the 1988 soft x-ray transient GS 2000+25. Using the known photometric period of 8.25836 hr (Chevalier and Ilovaisky 1993, A.Ap. 269, 301), we measure a velocity semiamplitude of 490 ± 22 km/s for the secondary star by cross-correlation with a K5 V ‘template’. This leads to a mass function of 4.2 ± 0.5 M_{⊙}, which—combined with the absence of eclipses and the observed ellipsoidal modulation (Callanan and Charles 1991, MNRAS 249, 573)—suggests a compact object mass of > 5.5 M_{⊙} (for inclinations < 70^{o}), establishing it firmly as a strong blackhole candidate. This is further supported by the Hα double-peak separation of 1200 km/s, implying an outer-disk velocity of 600 ± 30 km/s, a value that is typically 10–20 percent greater than the K velocity in other x-ray transients (e.g., Orosz and Bailyn 1994, IAUC 6103). QZ Vul thus has the second largest mass function known amongst the blackhole x-ray binaries, after V404 Cyg (Casares and Charles 1994, MNRAS 271, L5).”

1995 July 28 Daniel W. E. Green